OpenStudy (anonymous):

find the sum of ((x+11)/(x^2+x-132)) and ((x-12)/(x^2+23x+132)).

OpenStudy (mertsj):

\[\frac{x+11}{(x+12)(x-11)}+\frac{x-12}{(x+11)(x+12)}\]

OpenStudy (anonymous):

thanks for helping me.

OpenStudy (mertsj):

\[\frac{(x+11)(x+11)}{(x+12)(x-11)(x+11)}+\frac{(x-12)(x-11)}{(x+12)(x-11)(x+11)}\]

OpenStudy (mertsj):

\[\frac{x^2+22x+121+x^2-23x+132}{(x+12)(x-11)(x+11}\]

OpenStudy (mertsj):

\[\frac{2x^2-x+253}{(x+12)(x-11)(x+11)}\]

OpenStudy (anonymous):

thank you Mertsj! is that the final product?

OpenStudy (mertsj):

It's not a product. It's a sum.

OpenStudy (mertsj):

and you're welcome

OpenStudy (anonymous):

oh i'm sorry. the sum, yes. thank you so much!